# Thermodynamic nonequilibrium effects in three-dimensional high-speed compressible flows: Multiscale modeling and simulation via the discrete Boltzmann method

### v1

@@ IV Capability to capture higher-order thermodynamic nonequilibrium effects
The previous examples primarily demonstrate the model’s ability to capture large-scale flow structures, focusing on the evolution of macroscopic, conserved, and slow varying variables. In contrast, this section evaluates the model’s ability to represent mesoscopic, nonconserved, and fast varying variables. In this study, the intensity of TNEs, as defined by the analytical expressions in Eqs. (33)-(34) and Table 1, depends on macroscopic quantities, their gradients, relaxation times, and other factors. Thus, an intensity vector $\mathbf{S}_{\text{TNE}}=(\tau,\rho,T,\mathbf{u},\bm{\nabla}\rho,\bm{\nabla}T,\bm{\nabla}\mathbf{u},\bm{\Delta}_{2}^{*},\bm{\Delta}_{31}^{*})$, comprising nine elements, is defined to fully characterize the nonequilibrium manifestations.
@@
$\rho(x,y)=\frac{\rho_{L}+\rho_{R}}{2}-\frac{\rho_{L}-\rho_{R}}{2}\tanh\left(\frac{x-N_{x}\Delta x/2}{L_{\rho}}\right),$ (50)
$u(x,y)=-u_{0}\tanh\left(\frac{x-N_{x}\Delta x/2}{L_{u}}\right),$ (51)
@@ IV.1 Typical thermodynamic nonequilibrium quantities: Viscous stress
The initial values of other physical quantities are set as $\rho_{L}=2\rho_{R}=2,P_{L}=2,P_{R}=1,u_{y}=0,L_{\rho}=L_{u}=20$. By varying $\tau$ and $u_{0}$, three different cases of viscous stress are examined: (i) $\tau=10^{-4},u_{0}=0$; (ii) $\tau=10^{-4},u_{0}=0.5$; and (iii) $\tau=4\times 10^{-3},u_{0}=1.5$. Figure 8 shows the viscous stress distributions at $t=0.0025$, calculated using both the NS-level D3V30 model and the Burnett-level D3V55 model. The NS-level model satisfies seven kinetic moment relations (Eqs. (12)-(18)), whereas the Burnett-level model requires nine (Eqs. (12)-(20)). In Fig. 8, the green dotted line represents the first-order analytical solution, the red solid line the second-order analytical solution, and the blue hollow circles the DBM numerical solution.
@@
Subsequently, they extended the relative nonequilibrium strength to a TNE vector $\mathbf{R}_{\text{TNE}}=\left|\bm{\Delta}_{mn}^{*(j+1)}/\bm{\Delta}_{mn}^{*(j)}\right|$ , where $\bm{\Delta}_{mn}^{*(j+1)}$ and $\bm{\Delta}_{mn}^{*(j)}$ represent the $(j+1)$-th and $j$-th order TNEs, respectively. The deviation $\Delta=\Delta_{\text{DBM}}-\Delta_{\text{Exact}}$ between the numerical solutions from DBM simulations and their analytical solutions is used to assess the suitability of DBM for the current problem. The current-order DBM accurately describes the problem only when both $R_{\text{TNE}}$ and $\Delta$ are sufficiently small. Otherwise, a higher-order DBM model should be constructed to capture the higher-order TNEs Gan et al. (2022).
@@
(ii) The differences in macroscopic quantities between the high-order and low-order models are correlated with the relative strength of nonequilibrium. As the relative nonequilibrium strength increases (0.75, 0.035, and 1.14), the discrepancies between the models for each macroscopic quantity grow significantly. This suggests that second-order TNEs play a substantial role in system evolution, highlighting the necessity for higher-order physical models.
@@
To further assess the reliability of the D3V55 model in capturing stronger TNEs, $\tau$ was increased to $6\times 10^{-3}$ and $u_{0}$ was raised to 1.9. At this point, $R_{\text{TNE}}$ reached a value of 4.39 (at $x=503$). As shown in Fig. 10 (a), $\Delta_{2xx}^{*(1)}$ failed to accurately capture the peak region, while the DBM numerical solution matches $\Delta_{2xx}^{*(1)}+\Delta_{2xx}^{*(2)}$ well. This indicates that the second-order stress dominates in high nonequilibrium regions, particularly where velocity and temperature gradients vary drastically. The D3V55 model clearly exhibits a stronger ability to describe TNEs compared to the 2D Burnett-level DBM. Figure 10 (b) shows the local Knudsen number calculated from density, temperature, and pressure. Its distribution along the $x$-axis is similar, but not identical, to the distribution of viscous stress, indicating that different macroscopic quantities have distinct characteristic structures. As a typical nonequilibrium quantity, the distribution of viscous stress depends on the distributions of all macroscopic quantities and their nonlinear coupling. The maximum Knudsen numbers calculated from density and temperature are 0.20 and 0.25, respectively, while the Knudsen number calculated from pressure reaches 0.81, indicating that the flow has entered the transition regime.
@@
Further investigations were conducted to examine the effects of velocity and relaxation time on viscous stress. Figure 11 illustrates how the maximum viscous stress $\Delta_{2xx-\text{max}}^{*}$ varies with relaxation time $\tau$ for different initial velocities $u_{0}$. The red solid lines represent the results of both linear and nonlinear fitting. The figure clearly shows that as relaxation time and flow velocity increase, the strength of the TNEs also grows. Additionally, a critical flow velocity $u_{c}(=1.6)$ exists: when $u_{0}$ is below this threshold, the maximum viscous stress $\Delta_{2xx-\text{max}}^{*}$ increases linearly with relaxation time $\tau$, as described by the relation $\Delta_{2xx-\text{max}}^{*}=A_{1}+B_{1}\tau$. However, when $u_{0}$ exceeds this threshold, the maximum viscous stress increases nonlinearly with relaxation time, in accordance with the relation $\Delta_{2xx-\text{max}}^{*}=A_{2}+B_{2}\tau+C\tau^{2}$. This indicates that, under strong nonequilibrium conditions, the linear constitutive relations used in the NS fluid model break down, and the physical model must account for higher-order deviations $f^{j}(j\geq 2)$ from equilibrium in the distribution function.
@@
The interface widths $L_{\rho}$ and $L_{u}$ also affect the TNEs. With $u_{0}=1.6$ and $\tau=4\times 10^{-3}$, Figure 12 demonstrates that as the interface widths $L_{\rho}$ and $L_{u}$ increase, the maximum viscous stress $\Delta_{2xx-\text{max}}^{*}$ gradually decreases. This occurs because increasing the interface width reduces the macroscopic gradient, thereby suppressing the strength of the TNEs. The relationship between $\Delta_{2xx-\text{max}}^{*}$ and $L_{\rho}$ (or $L_{u}$) follows a negative power-law: $\Delta_{2xx-\text{max}}^{*}=A_{1}\cdot L_{\rho}^{B_{1}}$, where $A_{1}\approx 21.65$ and $B_{1}\approx-0.92$. The curve fitting has a correlation coefficient $R^{2}=0.98433$, indicating the goodness of the fit.
@@ IV.2 Typical thermodynamic nonequilibrium quantity: Heat flux
In case I, the initial temperature is uniform, resulting in a nearly zero first-order heat flux, $\Delta_{31x}^{*(1)}$, as the system evolves. The second-order heat flux, $\Delta_{31x}^{*(2)}$, is initially triggered by the density gradient, and is further excited by the temperature and velocity gradients, which lead to the emergence of the first-order heat flux. As shown in Fig. 13 I(a), when the second-order TNEs dominate, the heat flux computed by the D3V30 model deviates significantly from the analytical solution. In contrast, the D3V55 model’s numerical solution closely matches the second-order analytical solution, $\Delta_{31x}^{*(1)}+\Delta_{31x}^{*(2)}$, emphasizing the necessity of accounting for second-order TNEs.
@@
In case II, increasing the temperature gradient and velocity significantly enhances the amplitude of the heat flux, contributed by both the first-order and second-order TNEs. However, in this case, $R_{\text{TNE}}=0.08$, and the second-order nonequilibrium intensity remains relatively weak. Nonetheless, the D3V30 model fails to accurately describe the system’s heat flux, with both its amplitude and distribution deviating significantly from the analytical solution. As shown in Fig. 13 II(b), the D3V55 model more accurately captures the second-order TNEs, with its numerical solution closely matching the second-order analytical solution. The second-order quantity, $\Delta_{31x}^{*(2)}$, provides weak positive feedback to the overall nonequilibrium.
@@
Figure 14 presents a comparison of macroscopic quantities computed using the D3V55 model with those from the D3V30 model for case III at $t=0.00175$. The computed difference in macroscopic quantities for case III is $22\%$, indicating that, in strong nonequilibrium situations, the distribution of macroscopic quantities deviates from the continuous medium assumption, exhibiting significant discretization. In Burnett-level heat flux cases, the differences between the two models in each macroscopic quantity gradually increase with the accumulation of relative nonequilibrium strength, $R_{\text{TNE}}$, over time.
@@ V Conclusions, discussions and outlooks
3D high-speed compressible flows represent a complex interplay of nonlinear, nonequilibrium, and multiscale phenomena. These flows present formidable challenges to traditional fluid models, particularly in capturing the significant discrete effects and intricate thermodynamic nonequilibrium effects (TNEs) that emerge under extreme conditions. To address these challenges, this study develops a discrete Boltzmann modeling and simulation method, which fuses the insights of kinetic theory with mean-field theory. By using the Chapman-Enskog multiscale method, a set of kinetic moment relations $\bm{\Phi}=(\mathbf{M}_{0},\mathbf{M}_{1},\mathbf{M}_{20},\mathbf{M}_{2},\mathbf{M}_{31},\mathbf{M}_{3},\mathbf{M}_{42},\mathbf{M}_{4},\mathbf{M}_{53})$ for describing second-order TNEs are identified. These kinetic moment relations $\bm{\Phi}$ are crucial invariants that must be preserved in coarse-grained physical modeling, providing a fresh and deeper perspective for analyzing the TNE behaviors and their effects. A discrete Boltzmann model, employing 55 discrete velocities to discretize the 3D velocity space, is proposed for multiscale simulations of 3D supersonic flows. The nonlinear constitutive relations, as a by-product, are also derived, contributing to the theoretical foundation of macroscopic fluid modeling. Compared to the 2D case, the nonlinear constitutive relations in three dimensions are more complex due to the increased degrees of freedom, which lead to a greater variety of nonequilibrium driving forces, enhanced coupling between different non-equilibrium driving forces, and a significant increase in nonequilibrium components. Extensive numerical validations demonstrate the model’s accuracy across test cases ranging from 1D to 3D scenarios, and from subsonic to supersonic regimes. Particularly, the model effectively captures multiscale TNEs, including viscous stress and heat flux, near mesoscale structures.
@@
In principle, the extended Burnett equations (EBE) can also be used to study these mesoscale behaviors. However, deriving the full EBE is inherently difficult, and its high degree of nonlinearity and numerous nonlinear terms pose significant challenges for numerical simulations. These combined complexities render the approach of “first deriving the EBE, then numerically solving it” almost impractical. The DBM enables some of these studies, which become increasingly infeasible with higher Knudsen numbers, feasible. Moreover, compared to the kinetic macro modeling (KMM) approach of“first deriving the EBE, then solving it numerically”, the DBM offers new methodologies for analyzing the vast datasets generated from simulations. It particularly excels in answering the critical questions of “how to describe discrete states and effects”, and “ how to extract and analyze nonequilibrium states and effects” in a rigorous and efficient manner.
@@
Despite its strengths, the D3V55 DBM does have certain limitations. Despite its strengths, the D3V55 DBM has certain limitations. First, it currently accounts only for second-order TNEs, excluding higher-order effects that become crucial as the Mach number increases and compressibility strengthens. Second, the utilization of a single relaxation time simplifies modeling and computations, but fails to capture the varying rates at which different kinetic modes approach equilibrium, particularly when the disparities are prominent in complex or strongly nonequilibrium flows. Additionally, stability issues can also arise in some cases, especially at high Knudsen and Mach numbers.
@@
Future research will address these limitations in three key areas: (1) Incorporating higher-order TNEs, such as third-order effects, to improve accuracy under extreme nonequilibrium conditions; (2) Developing multi-relaxation-time higher-order models Yan and Shan (2024); Li et al. (2024) to better account for the varying timescales of different kinetic modes; and (3) Conducting von Neumann stability analyses Gan et al. (2008) to identify factors influencing model stability and propose strategies to enhance its robustness and reliability across diverse flow conditions. These advancements are expected to broaden the model’s applicability to aerospace engineering, energy extraction, security and defense, and other advanced technological fields.

### v2

@@ IV Capability to capture higher-order thermodynamic nonequilibrium effects
The previous examples primarily demonstrate the model’s ability to capture large-scale flow structures, focusing on the evolution of macroscopic, conserved, and slow varying variables. In contrast, this section evaluates the model’s ability to represent mesoscopic, nonconserved, and fast varying variables. In this study, the intensity of TNEs, as defined by the analytical expressions in Eqs. (35)-(36) and Table 1, depends on macroscopic quantities, their gradients, relaxation times, and other factors. Thus, an intensity vector $\mathbf{S}_{\text{TNE}}=(\tau,\rho,T,\mathbf{u},\bm{\nabla}\rho,\bm{\nabla}T,\bm{\nabla}\mathbf{u},\bm{\Delta}_{2}^{*},\bm{\Delta}_{3,1}^{*})$, comprising nine elements, is defined to fully characterize the nonequilibrium manifestations.
The nonequilibrium intensity vector offers a multi-perspective description of nonequilibrium states and behaviors. Its components consist of characteristic quantities and key quantities of interest. Characteristic quantities are defined by the system’s nonequilibrium driving mechanisms, while quantities of interest are determined by the system’s characteristics and are highlighted in both traditional computational fluid dynamics and the DBM framework. Additional components can be incorporated into $\mathbf{S}_{\text{TNE}}$ as necessary to maintain the vector’s completeness.
@@
$\rho(x,y,z)=\frac{\rho_{L}+\rho_{R}}{2}-\frac{\rho_{L}-\rho_{R}}{2}\tanh\left(\frac{x-N_{x}\Delta x/2}{L_{\rho}}\right),$ (52)
$u_{x}(x,y,z)=-u_{0}\tanh\left(\frac{x-N_{x}\Delta x/2}{L_{u}}\right),$ (53)
@@ IV.1 Typical thermodynamic nonequilibrium quantities: Viscous stress
The initial values of other physical quantities are set as $\rho_{L}=2\rho_{R}=2,P_{L}=2,P_{R}=1,u_{y}=0,u_{z}=0,L_{\rho}=L_{u}=20$. By varying $\tau$ and $u_{0}$, three different cases of viscous stress are examined: (i) $\tau=10^{-4},u_{0}=0$; (ii) $\tau=10^{-4},u_{0}=0.5$; and (iii) $\tau=4\times 10^{-3},u_{0}=1.5$. Figure 8 shows the viscous stress distributions at $t=0.0025$, calculated using both the NS-level D3V30 model and the Burnett-level D3V55 model. The NS-level model satisfies seven kinetic moment relations [Eqs. (12)-(18)], whereas the Burnett-level model requires nine [Eqs. (12)-(20)]. In Fig. 8, the green dotted line represents the first-order analytical solution, the red solid line the second-order analytical solution, and the blue hollow circles the DBM numerical solution.
@@
Subsequently, they extended the relative nonequilibrium strength to a TNE vector $\bm{R}_{\text{TNE}}=\left|\bm{\Delta}_{m,n}^{*(j+1)}/\bm{\Delta}_{m,n}^{*(j)}\right|$ , where $\bm{\Delta}_{m,n}^{*(j+1)}$ and $\bm{\Delta}_{m,n}^{*(j)}$ represent the $(j+1)$th and $j$th order TNEs, respectively. The deviation $\Delta=\Delta_{\text{DBM}}-\Delta_{\text{Exact}}$ between the numerical solutions from DBM simulations and their analytical solutions is used to assess the suitability of DBM for the current problem. The current-order DBM accurately describes the problem only when both $R_{\text{TNE}}$ and $\Delta$ are sufficiently small. Otherwise, a higher-order DBM model should be constructed to capture the higher-order TNEs Gan et al. (2022).
@@
(ii) The differences in macroscopic quantities between the high-order and low-order models are correlated with the relative strength of nonequilibrium. As the relative nonequilibrium strength increases (0.75, 0.035, and 1.14), the discrepancies between the models for each macroscopic quantity grow significantly. This suggests that second-order TNEs play a substantial role in system evolution, highlighting the necessity for higher-order physical models. Specifically, low-order nonequilibrium effects, represented by $\bm{\Delta}_{2}^{*}$ and $\bm{\Delta}_{3,1}^{*}$, function as constitutive relations and are embedded in the macroscopic equations, thereby influencing the evolution of macroscopic variables under different dissipation mechanisms. High-order nonequilibrium effects, represented by $\bm{\Delta}_{3}^{*}$ and $\bm{\Delta}_{4,2}^{*}$, serve as constitutive relations for $\bm{\Delta}_{2}^{*}$ and $\bm{\Delta}_{3,1}^{*}$, thereby governing the evolution equations of lower-order constitutive relationsGan et al. (2022).
@@
To further assess the reliability of the D3V55 model in capturing stronger TNEs, $\tau$ was increased to $6\times 10^{-3}$ and $u_{0}$ was raised to 1.9. At this point, $R_{\text{TNE}}$ reached a value of 4.39 (at $x=503$). As shown in Fig. 10 (a), $\Delta_{2xx}^{*(1)}$ failed to accurately capture the peak region, while the DBM numerical solution matches $\Delta_{2xx}^{*(1)}+\Delta_{2xx}^{*(2)}$ well. This indicates that the second-order stress dominates in high nonequilibrium regions, particularly where velocity and temperature gradients vary drastically. The D3V55 model clearly exhibits a stronger ability to describe TNEs compared to the 2D Burnett-level DBM. Figure 10 (b) shows the local Knudsen number calculated from density, temperature, and pressure. Its distribution along the $x$-axis is similar, but not identical, to the distribution of viscous stress, indicating that different macroscopic quantities have distinct characteristic structures. This is because (i) their dominant and dissipation mechanisms differ, and (ii) the Knudsen number, calculated from the gradient of conserved quantities, is a slowly varying and simpler variable, whereas nonequilibrium quantities are fast-varying, nonlinear combinations of macroscopic variables and their gradients, which more accurately reflect the system’s deviation from equilibrium and corresponding effects. As a typical nonequilibrium quantity, the distribution of viscous stress depends on the distributions of all macroscopic quantities and their nonlinear coupling. The maximum Knudsen numbers calculated from density and temperature are 0.20 and 0.25, respectively, while the Knudsen number calculated from pressure reaches 0.81, indicating that the flow has entered the transition regime.
@@
Further investigations were conducted to examine the effects of velocity and relaxation time on viscous stress. Figure 11 illustrates how the maximum viscous stress $\Delta_{2xx-\text{max}}^{*}$ varies with relaxation time $\tau$ for different initial velocities $u_{0}$. The red solid lines represent the results of both linear and nonlinear fitting. The figure clearly shows that as relaxation time and flow velocity increase, the strength of the TNEs also grows. Additionally, a critical flow velocity $u_{c}(=1.6)$ exists: when $u_{0}$ is below this threshold, the maximum viscous stress $\Delta_{2xx-\text{max}}^{*}$ increases linearly with relaxation time $\tau$, as described by the relation $\Delta_{2xx-\text{max}}^{*}=A_{1}+B_{1}\tau$. However, when $u_{0}$ exceeds this threshold, the maximum viscous stress increases nonlinearly with relaxation time, in accordance with the relation $\Delta_{2xx-\text{max}}^{*}=A_{2}+B_{2}\tau+C\tau^{2}$. This indicates that, under strong nonequilibrium conditions, the linear constitutive relations used in the NS fluid model break down, and the physical model must account for higher-order deviations $f^{j}(j\geq 2)$ from equilibrium in the distribution function. Similarly, the flow velocity has a comparable effect on other TNE quantities. In fact, the magnitude of the relative nonequilibrium strength determines the transition between linear and nonlinear behavior.
@@
The interface widths $L_{\rho}$ and $L_{u}$ also affect the TNEs. With $u_{0}=1.6$ and $\tau=4\times 10^{-3}$, Figure 12 demonstrates that as the interface widths $L_{\rho}$ and $L_{u}$ increase, the maximum viscous stress $\Delta_{2xx-\text{max}}^{*}$ gradually decreases. This occurs because increasing the interface width reduces the macroscopic gradient, thereby suppressing the strength of the TNEs. The relationship between $\Delta_{2xx-\text{max}}^{*}$ and $L_{\rho}$ (or $L_{u}$) follows a negative power-law: $\Delta_{2xx-\text{max}}^{*}=A_{1}\cdot L_{\rho}^{B_{1}}$, where $A_{1}\approx 21.65$ and $B_{1}\approx-0.92$. The curve fitting has a correlation coefficient $R^{2}=0.98$, indicating the goodness of the fit.
@@ IV.2 Typical thermodynamic nonequilibrium quantity: Heat flux
In case I, the initial temperature is uniform, resulting in a nearly zero first-order heat flux, $\Delta_{3,1x}^{*(1)}$, as the system evolves. The second-order heat flux, $\Delta_{3,1x}^{*(2)}$, is initially triggered by the density gradient, and is further excited by the temperature and velocity gradients, which lead to the emergence of the first-order heat flux. As shown in Fig. 13 I(a), when the second-order TNEs dominate, the heat flux computed by the D3V30 model deviates significantly from the analytical solution. In contrast, the D3V55 model’s numerical solution closely matches the second-order analytical solution, $\Delta_{3,1x}^{*(1)}+\Delta_{3,1x}^{*(2)}$, emphasizing the necessity of accounting for second-order TNEs.
@@
In case II, increasing the temperature gradient and velocity significantly enhances the amplitude of the heat flux, contributed by both the first-order and second-order TNEs. However, in this case, $R_{\text{TNE}}=0.08$, and the second-order nonequilibrium intensity remains relatively weak. Nonetheless, the D3V30 model fails to accurately describe the system’s heat flux, with both its amplitude and distribution deviating significantly from the analytical solution. As shown in Fig. 13 II(b), the D3V55 model more accurately captures the second-order TNEs, with its numerical solution closely matching the second-order analytical solution. The second-order quantity, $\Delta_{3,1x}^{*(2)}$, provides weak positive feedback to the overall nonequilibrium.
@@
Figure 14 presents a comparison of macroscopic quantities computed using the D3V55 model with those from the D3V30 model for case III at $t=0.00175$. The differences arise from variations in their ability to describe thermodynamic nonequilibrium effects. Some nonequilibrium quantities are directly embedded into the macroscopic fluid equations, determining the evolution and dissipation of macroscopic quantities. The computed difference in macroscopic quantities for case III is $22\%$, indicating that, in strong nonequilibrium situations, the distribution of macroscopic quantities deviates from the continuous medium assumption, exhibiting significant discretization. In Burnett-level heat flux cases, the differences between the two models in each macroscopic quantity gradually increase with the accumulation of relative nonequilibrium strength, $R_{\text{TNE}}$, over time. These differences, due to the physical model’s accuracy, will not diminish with improvements in numerical precision.
@@ V Conclusions, discussions and outlooks
Three-dimensional (3D) high-speed compressible flows represent a complex interplay of nonlinear, nonequilibrium, and multiscale phenomena. These flows present formidable challenges to traditional fluid models, particularly in capturing the significant discrete effects and intricate thermodynamic nonequilibrium effects (TNEs) that emerge under extreme conditions. To address these challenges, this study develops a discrete Boltzmann modeling and simulation method, which fuses the insights of kinetic theory with mean-field theory. By using the Chapman-Enskog multiscale method, a set of kinetic moment relations $\bm{\Phi}=(\mathbf{M}_{0},\mathbf{M}_{1},\mathbf{M}_{2,0},\mathbf{M}_{2},\mathbf{M}_{3,1},\mathbf{M}_{3},\mathbf{M}_{4,2},\mathbf{M}_{4},\mathbf{M}_{5,3})$ for describing second-order TNEs are identified. These kinetic moment relations $\bm{\Phi}$ are crucial invariants that must be preserved in coarse-grained physical modeling, providing a fresh and deeper perspective for analyzing the TNE behaviors and their effects. A discrete Boltzmann model, employing 55 discrete velocities to discretize the 3D velocity space, is proposed for multiscale simulations of 3D supersonic flows. The nonlinear constitutive relations, as a by-product, are also derived, contributing to the theoretical foundation of macroscopic fluid modeling. Compared to the 2D case, the nonlinear constitutive relations in three dimensions are more complex due to the increased degrees of freedom, which lead to a greater variety of nonequilibrium driving forces, enhanced coupling between different nonequilibrium driving forces, and a significant increase in nonequilibrium components. Extensive numerical validations demonstrate the model’s accuracy across test cases ranging from 1D to 3D scenarios, and from subsonic to supersonic regimes. Particularly, the model effectively captures multiscale TNEs, including viscous stress and heat flux, near mesoscale structures.
@@
In principle, the extended Burnett equations (EBE) can also be used to study these mesoscale behaviors. However, deriving the full EBE is inherently difficult, and its high degree of nonlinearity and numerous nonlinear terms pose significant challenges for numerical simulations. These combined complexities render the approach of “first deriving the EBE, then numerically solving it” almost impractical. The DBM enables some of these studies, which become increasingly infeasible with higher Knudsen numbers, feasible. Moreover, compared to the kinetic macro modeling (KMM) approach of “first deriving the EBE, then solving it numerically”, the DBM offers new methodologies for analyzing the vast datasets generated from simulations. It particularly excels in answering the critical questions of “how to describe discrete states and effects”, and “how to extract and analyze nonequilibrium states and effects” in a rigorous and efficient manner. In this study, we assess the model’s applicability by evaluating the self-consistency of the numerical and analytical solutions for nonequilibrium effects, as well as the relative nonequilibrium intensity. The numerical results (Fig. 10) show that when the Knudsen number computed from the density gradient reaches 0.2 (with Knudsen numbers of 0.25 and 0.81 calculated from the temperature and pressure gradients, respectively), the model proves physically effective and is suitable for studying nonequilibrium problems across continuum to transition flow regimes.
@@
Despite its strengths, the D3V55 DBM has certain limitations. First, it currently accounts only for second-order TNEs, excluding higher-order effects that become crucial as the Mach number increases and compressibility strengthens. Second, the utilization of a single relaxation time simplifies modeling and computations, but fails to capture the varying rates at which different kinetic modes approach equilibrium, particularly when the disparities are prominent in complex or strongly nonequilibrium flows. Additionally, stability issues can also arise in some cases, especially at high Knudsen and Mach numbers.
@@
Future research will address these limitations in three key areas: (1) Incorporating higher-order TNEs, such as third-order effects, to improve accuracy under extreme nonequilibrium conditions; (2) To better account for the varying timescales of different kinetic modes, researchers have developed multi-relaxation-time higher-order models Yan and Shan (2024); Li et al. (2024); and (3) Conducting von Neumann stability analyses Gan et al. (2008) to identify factors influencing model stability and propose strategies to enhance its robustness and reliability across diverse flow conditions. These advancements are expected to broaden the model’s applicability to aerospace engineering, energy extraction, security and defense, and other advanced technological fields.